# MI-NeRF: Learning a Single Face NeRF from Multiple Identities

### v1

@@ Abstract
Keywords:
Face Representation Multiple Identities Neural Radiance Fields
@@ 1 Introduction
Capturing the 4D dynamics and appearance of non-rigid motion of humans has long been a challenge for both computer vision and graphics. This task has broad applications, ranging from AR/VR and video games to virtual communication and the movie industry, all of which require the creation of photorealistic videos of the human face. Earlier approaches relied on 3D morphable models (3DMM) [17, 16, 51], while later methods have turned to generative adversarial networks (GANs) [26, 44, 43, 58]. GANs learn representations of facial dynamics from large datasets, containing video clips from multiple identities. To disentangle latent factors of variation, such as identity and expression, some works have imposed multilinear structures [48, 61, 18], building on the ideas of TensorFaces [55]. Despite their success, most GANs operate in the 2D image space and they do not model the 3D face geometry.
Figure 1: Our proposed multi-identity NeRF (MI-NeRF) learns a single network to model complex non-rigid human face motion of multiple identities from just monocular videos. In contrast, standard approaches train multiple per-identity NeRFs.
@@
Neural radiance fields (NeRF) have recently demonstrated photorealistic 3D modeling of both static [37, 4, 5, 34] and dynamic scenes [45, 33, 32, 39, 15, 40, 65], making them a popular choice for modeling human faces from monocular videos [15, 39, 40, 3]. Approaches that leverage a 3DMM prior and condition on expression parameters enable control of facial expressions, for applications such as expression transfer [15] and lip syncing [8]. Despite their high-quality results, NeRFs require expensive per-scene or per-identity optimization. Recent works [72, 13, 59, 46] propose techniques to reduce the training and rendering times. However, increasing the number of identities to hundreds can be expensive, since we would need one model per identity with corresponding learnable parameters. A few works learn more generic representations [62, 9, 52, 67, 29, 38, 25], but they require static settings and/or multiple input views during training.
@@
In this work, we propose MI-NeRF (multi-identity NeRF), a novel method that learns a single dynamic NeRF from monocular talking face videos of multiple identities. Using only a single network, it learns to model complex non-rigid human face motion, while disentangling identity and non-identity specific information. At the epicenter of our approach lies a multiplicative module that approximates the non-linear interactions between latent factors of variation, inspired by ideas that go back to TensorFaces [55]. This module learns a non-linear mapping of identity codes and facial expressions, based on the Hadamard product. To the best of our knowledge, this is the first method that learns a single unified face NeRF from monocular videos of multiple identities.
@@ 2 Related Work
Human Portrait Video Synthesis. Earlier approaches for video synthesis and editing of human faces are based on 3DMMs [17, 16, 51]. A 3DMM [6] is a parametric model that represents a face as a linear combination of principle axes for shape, texture, and expression, learned by principal component analysis (PCA). GAN-based networks have been later proposed for video synthesis [26, 49, 44], as well as for audio-driven talking faces [43, 70, 58]. GANs are trained on large datasets with video clips from multiple identities, learning diverse facial expressions and lip movements. However, they operate in a low resolution 2D image space and they cannot model the 3D face geometry. NeRFs [37] have recently become very popular, since they can represent the 3D face geometry and appearance, and generate photorealistic videos [15, 23, 39, 40]. Subsequent works [72, 13, 46] propose techniques to reduce the training and inference time.
@@
Multilinear Factor Analysis of Faces. Factors of variation, such as identity, expression, and illumination, affect the appearance of a human face in a portrait video. Disentangling those factors is challenging. Techniques like PCA can only find a single mode of variation [54]. TensorFaces [55] is an early approach that approximates different modes of variation using a multilinear tensor decomposition. Inspired by this, several works have proposed to learn multiplicative interactions to disentangle latent factors of variation [56, 50, 60]. Multilinear latent conditioning has also been proved beneficial for GANs and VAEs, in order to disentangle and edit face attributes [48, 61, 18, 11]. In this work, we propose a multiplicative module that conditions a NeRF and approximates the non-linear interactions between identity and non-identity specific information.
@@
Neural Radiance Fields. Implicit neural representations for modeling 3D scenes have recently gained a lot of attention. In particular, NeRFs [37, 4, 5, 34] have shown photorealistic novel view synthesis of complex scenes. They represent a static scene as a continuous 5D function, using a multilayer perceptron (MLP) that maps each 5D coordinate (3D spatial location and 2D viewing direction) to an RGB color and volume density. However, NeRFs require expensive per-scene or per-identity optimization. A few works have proposed to learn generic representations [62, 9, 52, 67], but these require static settings and multiple views as input. In contrast, in this work, we are interested in dynamic human faces, captured from monocular videos.
@@
Dynamic Neural Radiance Fields for Human Faces. Several works have extended NeRFs to dynamic scenes [45, 33, 32, 39, 15, 40, 65]. They usually map the sampled 3D points from an observation space to a canonical space, in order to learn a time-invariant scene representation. Some additionally learn time-variant latent codes [15, 32, 8]. Particularly challenging is to capture the 4D dynamics and appearance of the non-rigid deformations of the human face from monocular videos. Conditioning a NeRF on 3DMM expression parameters, related works enable explicit and meaningful control of the synthesized subject [15, 23, 3, 2, 8, 72, 13]. Although these approaches can produce HD quality results, they are identity-specific and usually require long (more than a few seconds) videos for training. Only a limited prior work has aimed to train a generic NeRF for human faces [47, 25, 71]. However, all of these models require multiple views for training. In contrast, we propose a simple architecture that is capable of learning multiple identities from monocular videos of arbitrary length captured in the wild.
@@ 3.1 Conditional Input
Head Pose and Expression. For each video frame of an identity, we fit a 3DMM and extract the corresponding head pose $\bm{P}\in\mathbb{R}^{4\times 4}$ and expression parameters $\bm{e}\in\mathbb{R}^{\text{79}}$. We follow an optimization-based method, that minimizes an objective function with photo-consistency and landmark terms [22]. We use the learned axes from [22], based on the Basel Face Model [42] for shape and texture and the FaceWarehouse [7] for expression. The extracted head pose $\bm{P}=\left[\bm{R};\textbf{t}\right]$ is used to transform the sampled 3D points to the canonical space before shooting the rays, where $\bm{R}\in\mathbb{R}^{3\times 3}$ and $\textbf{t}~{}\in~{}\mathbb{R}^{3\times 1}$ are the rotation and translation matrices correspondingly.
Learned Identity and Latent Codes. In addition to the expression vectors, the dynamic NeRF is conditioned on learned identity and latent codes. Both are randomly initialized embeddings that are learned during training. We use one identity code $\bm{i}\in\mathbb{R}^{\text{79}}$ per video  [^1] In our preliminary experiments, we found that defining the identity vector with the same dimension as the expression vector was sufficient., in order to capture time-invariant information. These codes appear to mainly capture the identity, and thus we call them identity codes. We also learn one latent code $\bm{l}\in\mathbb{R}^{\text{32}}$ per frame per video, in order to capture time-varying information. These latent codes memorize very small per-frame variations, such as appearance and illumination, that are independent of facial expressions, but necessary to reconstruct them in the output videos [15, 8].
@@ 3.2 Proposed Modules
We propose to learn multiplicative interactions between facial expressions and identity codes using the Hadamard product. Earlier works have used multilinear tensor decomposition to disentangle latent factors of variation of the human face [55, 61, 18]. Inspired by this, we learn a non-linear mapping that disentangles identity and non-identity specific information. This mapping is learned by a single multiplicative module $M$ with shared weights for all identities. We also introduce a variation of this module, $H$, that captures high-degree interactions. Please see the appendix for the detailed derivation.
3.2.1 Multiplicative Interaction Module.
@@
where * denotes the Hadamard (element-wise) product that correlates $\bm{e}$ and $\bm{i}$, $\bm{U}_{1}\in\mathbb{R}^{k\times d}$, $\bm{U}_{2}\in\mathbb{R}^{k\times d}$, $\bm{C}\in\mathbb{R}^{d\times k}$, $\bm{W}_{2}\in\mathbb{R}^{d\times d}$, $\bm{W}_{3}\in\mathbb{R}^{d\times d}$ are learnable parameters, and $d=79$ for our case. We chose $k<d$ to get low rank matrices, with fewer parameters. We experimentally found that this mapping $M$ with $k=8$ leads to the best disentanglement between identity and expression with the least number of parameters (see ablation study in Sec. 4.2 for more details). Prop. 1 verifies the multiplicative interactions learned; its proof exists in the appendix in Sec A.1.
Proposition 1
@@
3.2.2 High-degree Interaction Module.
@@
for $n=2,\dots,N$, with $\bm{x}_{1}=\bm{U}_{(1,1)}\bm{e}+\bm{U}_{(1,2)}\bm{i}$. The parameters $\bm{U}_{(n,1)}\in\mathbb{R}^{k\times d}$, $\bm{U}_{(n,2)}\in\mathbb{R}^{k\times d}$, $\bm{C}\in\mathbb{R}^{o\times k}$ are learnable for $n=1,\dots,N$. In practice, we choose $d=k=o=79$ for our experiments. Using our proposed module $H$ with $N=2$ leads to similar performance with our proposed module $M$, and better performance compared to increasing $N$ to larger values (see ablation study in Sec. 4.2). However, we believe that capturing higher-degree interactions might be beneficial in other cases. Propositions 2 and 3 in Sec. A.2 in the appendix demonstrate the interactions learned in this case.
@@ 3.3 Dynamic NeRF
To model the dynamics of human faces, we learn a single dynamic NeRF for all input identities. For each video frame of a subject, we first segment the head using an automatic parsing method [30], similarly with [15, 23, 8]. Then, we learn an implicit representation $F_{\Theta}$ of the identities using an MLP. Given an identity $\bm{i}$ at a specific video frame, shown from a particular viewpoint and with a particular facial expression, we first march camera rays through the scene and sample 3D points on these rays. For a 3D point location $\bm{x}$, a viewing direction $\bm{v}$, the estimated expression vector $\bm{e}$, and a learned latent vector $\bm{l}$, $F_{\Theta}$ predicts the RGB color $\bm{c}$ and density $\sigma$ of the point:
$F_{\Theta}:(M(\bm{e},\bm{i}),\bm{l},\bm{x},\bm{v})\longrightarrow(\bm{c},\sigma)\;,$ (3)
where $M(\bm{e},\bm{i})$ can also be replaced with $H(\bm{e},\bm{i})$ (see Sec. 3.2). Given the predicted color $\bm{c}$ and density $\sigma$ for every point on each ray, we can produce the final video frame applying volumetric rendering [37]. For each camera ray $\bm{r}(t)=\bm{o}+t\bm{v}$ with camera center $\bm{o}$ and viewing direction $\bm{v}$, the color $C$ of the corresponding pixel can be computed by accumulating the predicted colors and densities of the sampled points along the ray:
@@
Implementation Details. We use an MLP of 8 linear layers with a hidden size of 256 and ReLU activations, with branches for $\bm{c}$ and $\sigma$, positional encodings for $\bm{x}$ and $\bm{v}$ of 10 and 4 frequencies respectively, and Adam optimizer [27] with a learning rate of $5\times 10^{-4}$ that decays exponentially to $5\times 10^{-5}$ (see also appendix).
@@ 3.4 Personalization
Trained on multiple identities simultaneously, MI-NeRF learns a large variety of facial expressions from diverse human faces and can synthesize videos of any training identity. To enhance the visual quality for a particular seen subject, i.e., to better capture their facial details, such as wrinkles, we can further improve the output appearance using a short video of this subject. We call this procedure “personalization”. More specifically, we fine-tune the network with a small learning rate ($10^{-5}$) for only a few iterations, keeping the weights of the multiplicative module frozen. This idea can also be used to adapt MI-NeRF to an unseen identity (that is not part of the initial training set). Given only a few frames of an unseen subject, we can fine-tune our network to learn their identity and latent codes. Then, we can synthesize high-quality videos of them, given novel expressions as input.

### v2

@@ 1 Introduction
Capturing the 4D dynamics and appearance of non-rigid motion of humans has long been a challenge for both computer vision and graphics. This task has broad applications, ranging from AR/VR and video games to virtual communication and the movie industry, all of which require the creation of photorealistic videos of the human face. Earlier approaches relied on 3D morphable models (3DMM) (Garrido et al., 2015; 2014; Thies et al., 2016), while later methods have turned to generative adversarial networks (GANs) (Kim et al., 2018; Pumarola et al., 2020; Prajwal et al., 2020; Vougioukas et al., 2020). GANs learn representations of facial dynamics from large datasets, containing video clips from multiple identities. To disentangle latent factors of variation, such as identity and expression, some works have imposed multilinear structures (Sahasrabudhe et al., 2019; Wang et al., 2019; Georgopoulos et al., 2020), building on the ideas of TensorFaces (Vasilescu & Terzopoulos, 2002). Despite their success, most GANs operate in the 2D image space and they do not model the 3D face geometry.
@@
Neural radiance fields (NeRF) have recently demonstrated photorealistic 3D modeling of both static (Mildenhall et al., 2020; Barron et al., 2021; 2022; Lindell et al., 2022) and dynamic scenes (Pumarola et al., 2021; Li et al., 2021; 2022; Park et al., 2021a; Gafni et al., 2020; Park et al., 2021b; Weng et al., 2022), making them a popular choice for modeling human faces from monocular videos (Gafni et al., 2020; Park et al., 2021a; b; Athar et al., 2022). Approaches that leverage a 3DMM prior and condition on expression parameters enable control of facial expressions, for applications such as expression transfer (Gafni et al., 2020) and lip syncing (Chatziagapi et al., 2023). Despite their high-quality results, NeRFs require expensive per-scene or per-identity optimization. Recent works (Zielonka et al., 2023; Duan et al., 2023; Wang et al., 2024; Qian et al., 2023) propose techniques to reduce the training and rendering times. However, increasing the number of identities to hundreds can be expensive, since we would need one model per identity with corresponding learnable parameters. A few works learn more generic representations (Wang et al., 2021a; Chen et al., 2021; Trevithick & Yang, 2021; Yu et al., 2021; Kwon et al., 2021; Mu et al., 2023; Hong et al., 2022), but they require static settings and/or multiple input views during training.
@@
In this work, we propose MI-NeRF (multi-identity NeRF), a novel method that learns a single dynamic NeRF from monocular talking face videos of multiple identities. Using only a single network, it learns to model complex non-rigid human face motion, while disentangling identity and non-identity specific information. At the epicenter of our approach lies a multiplicative module that approximates the non-linear interactions between latent factors of variation, inspired by ideas that go back to TensorFaces (Vasilescu & Terzopoulos, 2002). This module learns a non-linear mapping of identity codes and facial expressions, based on the Hadamard product. To the best of our knowledge, this is the first method that learns a single unified face NeRF from monocular videos of multiple identities.
@@ 2 Related Work
Human Portrait Video Synthesis. Earlier approaches for video synthesis and editing of human faces are based on 3DMMs (Garrido et al., 2015; 2014; Thies et al., 2016). A 3DMM (Blanz & Vetter, 1999) is a parametric model that represents a face as a linear combination of principle axes for shape, texture, and expression, learned by principal component analysis (PCA). GAN-based networks have been later proposed for video synthesis (Kim et al., 2018; Siarohin et al., 2019; Pumarola et al., 2020), as well as for audio-driven talking faces (Prajwal et al., 2020; Zhou et al., 2021; Vougioukas et al., 2020). GANs are trained on large datasets with video clips from multiple identities, learning diverse facial expressions and lip movements. However, they operate in a low resolution 2D image space and they cannot model the 3D face geometry. NeRFs (Mildenhall et al., 2020) have recently become very popular, since they can represent the 3D face geometry and appearance, and generate photorealistic videos (Gafni et al., 2020; Guo et al., 2021; Park et al., 2021a; b). Subsequent works (Zielonka et al., 2023; Duan et al., 2023; Qian et al., 2023) propose techniques to reduce the training and inference time.
@@
Multilinear Factor Analysis of Faces. Factors of variation, such as identity, expression, and illumination, affect the appearance of a human face in a portrait video. Disentangling those factors is challenging. Techniques like PCA can only find a single mode of variation (Turk & Pentland, 1991). TensorFaces (Vasilescu & Terzopoulos, 2002) is an early approach that approximates different modes of variation using a multilinear tensor decomposition. Inspired by this, several works have proposed to learn multiplicative interactions to disentangle latent factors of variation (Vlasic et al., 2006; Tang et al., 2013; Wang et al., 2017). Multilinear latent conditioning has also been proved beneficial for GANs and VAEs, in order to disentangle and edit face attributes (Sahasrabudhe et al., 2019; Wang et al., 2019; Georgopoulos et al., 2020; Chrysos et al., 2021). In this work, we propose a multiplicative module that conditions a NeRF and approximates the non-linear interactions between identity and non-identity specific information.
@@
Neural Radiance Fields. Implicit neural representations for modeling 3D scenes have recently gained a lot of attention. In particular, NeRFs (Mildenhall et al., 2020; Barron et al., 2021; 2022; Lindell et al., 2022) have shown photorealistic novel view synthesis of complex scenes. They represent a static scene as a continuous 5D function, using a multilayer perceptron (MLP) that maps each 5D coordinate (3D spatial location and 2D viewing direction) to an RGB color and volume density. However, NeRFs require expensive per-scene or per-identity optimization. A few works have proposed to learn generic representations (Wang et al., 2021a; Chen et al., 2021; Trevithick & Yang, 2021; Yu et al., 2021), but these require static settings and multiple views as input. In contrast, in this work, we are interested in dynamic human faces, captured from monocular videos.
@@
Dynamic Neural Radiance Fields for Human Faces. Several works have extended NeRFs to dynamic scenes (Pumarola et al., 2021; Li et al., 2021; 2022; Park et al., 2021a; Gafni et al., 2020; Park et al., 2021b; Weng et al., 2022). They usually map the sampled 3D points from an observation space to a canonical space, in order to learn a time-invariant scene representation. Some additionally learn time-variant latent codes (Gafni et al., 2020; Li et al., 2022; Chatziagapi et al., 2023). Particularly challenging is to capture the 4D dynamics and appearance of the non-rigid deformations of the human face from monocular videos. Conditioning a NeRF on 3DMM expression parameters, related works enable explicit and meaningful control of the synthesized subject (Gafni et al., 2020; Guo et al., 2021; Athar et al., 2022; 2021; Chatziagapi et al., 2023; Zielonka et al., 2023; Duan et al., 2023). Although these approaches can produce HD quality results, they are identity-specific and usually require long (more than a few seconds) videos for training. Only a limited prior work has aimed to train a generic NeRF for human faces (Raj et al., 2021; Hong et al., 2022; Zhuang et al., 2022). However, all of these models require multiple views for training. In contrast, we propose a simple architecture that is capable of learning multiple identities from monocular videos of arbitrary length captured in the wild.
@@ 3.1 Conditional Input
Head Pose and Expression. For each video frame of an identity, we fit a 3DMM and extract the corresponding head pose $\bm{P}\in\mathbb{R}^{4\times 4}$ and expression parameters $\bm{e}\in\mathbb{R}^{\text{79}}$. We follow an optimization-based method, that minimizes an objective function with photo-consistency and landmark terms (Guo et al., 2018b). We use the learned axes from Guo et al. (2018b), based on the Basel Face Model (Paysan et al., 2009) for shape and texture and the FaceWarehouse (Cao et al., 2013) for expression. The extracted head pose $\bm{P}=\left[\bm{R};\textbf{t}\right]$ is used to transform the sampled 3D points to the canonical space before shooting the rays, where $\bm{R}\in\mathbb{R}^{3\times 3}$ and $\textbf{t}~{}\in~{}\mathbb{R}^{3\times 1}$ are the rotation and translation matrices correspondingly.
Learned Identity and Latent Codes. In addition to the expression vectors, the dynamic NeRF is conditioned on learned identity and latent codes. Both are randomly initialized embeddings that are learned during training. We use one identity code $\bm{i}\in\mathbb{R}^{\text{79}}$ per video  [^1] In our preliminary experiments, we found that defining the identity vector with the same dimension as the expression vector was sufficient., in order to capture time-invariant information. These codes appear to mainly capture the identity, and thus we call them identity codes. We also learn one latent code $\bm{l}\in\mathbb{R}^{\text{32}}$ per frame per video, in order to capture time-varying information. These latent codes memorize very small per-frame variations, such as appearance and illumination, that are independent of facial expressions, but necessary to reconstruct them in the output videos (Gafni et al., 2020; Chatziagapi et al., 2023).
@@ 3.2 Proposed Modules
We propose to learn multiplicative interactions between facial expressions and identity codes using the Hadamard product. Earlier works have used multilinear tensor decomposition to disentangle latent factors of variation of the human face (Vasilescu & Terzopoulos, 2002; Wang et al., 2019; Georgopoulos et al., 2020). Inspired by this, we learn a non-linear mapping that disentangles identity and non-identity specific information. This mapping is learned by a single multiplicative module $M$ with shared weights for all identities. We also introduce a variation of this module, $H$, that captures high-degree interactions. Please see the appendix for the detailed derivation.
3.2.1 Multiplicative Interaction Module
@@
where * denotes the Hadamard (element-wise) product that correlates $\bm{e}$ and $\bm{i}$, $\bm{U}_{1}\in\mathbb{R}^{k\times d}$, $\bm{U}_{2}\in\mathbb{R}^{k\times d}$, $\bm{C}\in\mathbb{R}^{d\times k}$, $\bm{W}_{2}\in\mathbb{R}^{d\times d}$, $\bm{W}_{3}\in\mathbb{R}^{d\times d}$ are learnable parameters, and $d=79$ for our case. We chose $k<d$ to get low rank matrices, with fewer parameters. We experimentally found that this mapping $M$ with $k=8$ leads to the best disentanglement between identity and expression with the least number of parameters (see ablation study in Sec. 4.2 for more details). Prop. 1 verifies the multiplicative interactions learned; its proof exists in Sec. A.1.
Proposition 1.
@@
3.2.2 High-degree Interaction Module
@@
for $n=2,\dots,N$, with $\bm{x}_{1}=\bm{U}_{(1,1)}\bm{e}+\bm{U}_{(1,2)}\bm{i}$. The parameters $\bm{U}_{(n,1)}\in\mathbb{R}^{k\times d}$, $\bm{U}_{(n,2)}\in\mathbb{R}^{k\times d}$, $\bm{C}\in\mathbb{R}^{o\times k}$ are learnable for $n=1,\dots,N$. In practice, we choose $d=k=o=79$ for our experiments. Using our proposed module $H$ with $N=2$ leads to similar performance with our proposed module $M$, and better performance compared to increasing $N$ to larger values (see ablation study in Sec. 4.2). However, we believe that capturing higher-degree interactions might be beneficial in other cases. Prop. 2 and Prop. 3 in Sec. A.2 demonstrate the interactions learned in this case.
@@ 3.3 Dynamic NeRF
To model the dynamics of human faces, we learn a single dynamic NeRF for all input identities. For each video frame of a subject, we first segment the head using an automatic parsing method (Lee et al., 2020), similarly with Gafni et al. (2020); Guo et al. (2021); Chatziagapi et al. (2023). Then, we learn an implicit representation $F_{\Theta}$ of the identities using an MLP. Given an identity $\bm{i}$ at a specific video frame, shown from a particular viewpoint and with a particular facial expression, we first march camera rays through the scene and sample 3D points on these rays. For a 3D point location $\bm{x}$, a viewing direction $\bm{v}$, the estimated expression vector $\bm{e}$, and a learned latent vector $\bm{l}$, $F_{\Theta}$ predicts the RGB color $\bm{c}$ and density $\sigma$ of the point:
$F_{\Theta}:(M(\bm{e},\bm{i}),\bm{l},\bm{x},\bm{v})\longrightarrow(\bm{c},\sigma)\;.$ (3)
where $M(\bm{e},\bm{i})$ can also be replaced with $H(\bm{e},\bm{i})$ (see Sec. 3.2). Given the predicted color $\bm{c}$ and density $\sigma$ for every point on each ray, we can produce the final video frame applying volumetric rendering (Mildenhall et al., 2020). For each camera ray $\bm{r}(t)=\bm{o}+t\bm{v}$ with camera center $\bm{o}$ and viewing direction $\bm{v}$, the color $C$ of the corresponding pixel can be computed by accumulating the predicted colors and densities of the sampled points along the ray:
@@
Similarly to NeRF (Mildenhall et al., 2020), we follow a hierarchical sampling strategy, optimizing a coarse and a fine model. During training, we minimize the following objective function:
@@
Implementation Details. We use an MLP of 8 linear layers with a hidden size of 256 and ReLU activations, with branches for $\bm{c}$ and $\sigma$, positional encodings for $\bm{x}$ and $\bm{v}$ of 10 and 4 frequencies respectively, and Adam optimizer (Kingma & Ba, 2014) with a learning rate of $5\times 10^{-4}$ that decays exponentially to $5\times 10^{-5}$ (see appendix E for more details).
@@ 3.4 Personalization
Trained on multiple identities simultaneously, MI-NeRF learns a large variety of facial expressions from diverse human faces and can synthesize videos of any training identity. To enhance the visual quality for a particular seen subject, i.e. to better capture their facial details, such as wrinkles, we can further improve the output appearance using a short video of this subject. We call this procedure “personalization”. More specifically, we fine-tune the network with a small learning rate ($10^{-5}$) for only a few iterations, keeping the weights of the multiplicative module frozen. This idea can also be used to adapt MI-NeRF to an unseen identity (that is not part of the initial training set). Given only a few frames of an unseen subject, we can fine-tune our network to learn their identity and latent codes. Then, we can synthesize high-quality videos of them, given novel expressions as input.